Also the slope of
5x-y=-7

Answer by rahman(247)

5x-y=-7 -----------------> 5x+7=y or y = 5x + 7
So, the equation (5x - y = - 7) can modified to y = 5x + 7. And we just compare with y = mx + c; where m is the slope.
Thus, we get the slope of 5x - y = -7 is 5.
OK